OpenStudy (anonymous):

Find the general solution of \[cos x + \sqrt{3}sin x = \sqrt{2}\]

OpenStudy (anonymous):

\[\frac{1}{2}\cos x + \frac{\sqrt{3}}{2}\sin x = \frac{\sqrt{2}}{2}\]\[\cos \frac{\pi}{3}\times\cos x + \sin \frac{\pi}{3}\times\sin x = \frac{\sqrt{2}}{2}\]

OpenStudy (anonymous):

\[sin (x + pie/6) = \frac{1}{\sqrt{2}}\]

OpenStudy (anonymous):

be carefull\[\cos(x + \frac{\pi}{3}) = \cos \frac{\pi}{4} \]

OpenStudy (anonymous):

sorry i must be carefull too :)\[\cos(x - \frac{\pi}{3}) = \cos \frac{\pi}{4}\]

OpenStudy (anonymous):

Yes...i got it thx.....

OpenStudy (anonymous):

urs is right too